MEO Foods has made cat food for over 20 years. The company currently has a debt-equity ratio of 25%, borrows at 10% interest rate, and is in the 40% tax bracket. Its shareholders require an 18% return.

MEO is planning to expand cat food production capacity. The equipment to be purchased would last three years and generate the following unlevered cash flows(UCF):

Year 0: -$15 million.

Year 1: $5 million.

Year 2: $8 million.

Year 3: $10 million

Year 4: $0.

MEO has also arranged a $6 million debt issue to partially finance the expansion. Under the loan, the company would pay 10% annually on the outstanding balance. The firm would also make year-end principal payments of $2 million per year, completely retiring the issue at the end of the third year.

Ignoring the costs of financial distress and issue costs, what is the APV of the expansion plans? Please show work on how to reach the answer (given below).

Answer: $1.78 million
Answer #1
0 1 2 3 4
UCF -15.00 5.00 8.00 10.00 0.00
PVIF at 16.96% 1 0.8550 0.7310 0.6250 0.5344
PV at 16.96% -15.00 4.28 5.85 6.25 0.00
Value of unlevered operations 1.37
Beginning balance of debt 6.00 4.00 2.00 0.00
Interest at 10% 0.60 0.40 0.20 0.00
Tax shield at 40% 0.24 0.16 0.08 0
PVIF at 10% 0.9091 0.8264 0.7513 0.6830
PV at 10% 0.22 0.13 0.06 0.00
Value of tax shield on interest 0.41
APV of the expansion plan 1.78
Calculation for unlevered cost of equity:
rsl = rsu+(rsu-rd)*(1-t)*D/E
0.18 = rsu+(rsu-0.10)*0.60*0.25
0.18 = rsu+0.15*rsu-0.015
0.195 = 1.15*rsu
rsu = 0.195/1.15 = 16.96%